OpenStudy (knowledge):

integrate sqrt(1-x^2) Let x=cos theta

OpenStudy (anonymous):

ok let \(x=\cos(\theta)\) then \(\sqrt{1-x^2}=\sin(\theta)\) and \(dx=-\sin(\theta)\) your integral is \[-\int \sin^2(\theta)d\theta\]

OpenStudy (knowledge):

alright i am past that

OpenStudy (anonymous):

where you at?

OpenStudy (knowledge):

i have already integrated with respect to theta and have trouble substituting in

OpenStudy (anonymous):

ooh ok did you get to \[-\int\sin^2(\theta)=\frac{1}{2}(\sin(x)\cos(x)-1)\]?

OpenStudy (knowledge):

so i have (sin(2 theta)/4)-theta/2

OpenStudy (anonymous):

ok we can work with that one too, makes no difference

OpenStudy (knowledge):

how do i substitute in for theta

OpenStudy (knowledge):

sin(2theta)=2 sin theta cos theta=xsqrt(1-x^2)

OpenStudy (knowledge):

but what about theta

OpenStudy (anonymous):

\[\frac{x}{2}+\frac{1}{4}\sin(2\theta)\] want to work with that one?

OpenStudy (anonymous):

oooh you really did everything!

OpenStudy (anonymous):

you got the \(x\sqrt{1-x^2}\) part, that is the hard part \(\theta\) is the easy part \[x=\cos(\theta)\\ \theta=\cos^{-1}(x)\]

OpenStudy (knowledge):

but then many times i see arcsin(x) being used for theta instead of arccos(x)

OpenStudy (anonymous):

that is because you chose the substitution \(x=\cos(\theta)\) which makes \(\theta =\arccos(x)\)

OpenStudy (anonymous):

you could have chosen \(x=\sin(\theta)\) instead and done the integral that way

OpenStudy (anonymous):

\[x=\sin(\theta), dx=\cos(\theta)d\theta, \sqrt{1-x^2}=\cos(\theta)\] your integral is \[\int \cos^2(\theta)d\theta\]

OpenStudy (anonymous):

in this case you would use \[\cos^2(\theta)=\frac{1}{2}(1+\cos(2\theta))\]

OpenStudy (anonymous):

everything works out exactly the same and don't forget the the derivative of \(\arcsin(x)=\frac{1}{\sqrt{1-x^2}}\) while the derivative of \(\arccos(x)=-\frac{1}{\sqrt{1-x^2}}\) so it all works out exactly the same

OpenStudy (knowledge):

thanks

OpenStudy (anonymous):

yw